OpenStudy (anonymous):

Simplify the following expression. x^-1

hero (hero):

@faixy , you've been reported twice for giving answers without explanation just so you know.

OpenStudy (anonymous):

The easiest way to see this is to use the commutative property, then gather all the like factors into numbers or exponential expressions.\[4x^3y(-2x^2y)=4*3*-2*2x^3x^2yy=-48x^5y^2\]

OpenStudy (anonymous):

Wait a minute... the problem changed while I was working on it...! Your second problem posed is answered directly from the rules about exponents\[a^{-n}=\frac{1}{a^n}\implies x^{-1}=\frac{1}{x}\]